OpenStudy (johnkim2330):

What is the least common denominator of the fraction 3 x over x plus 1 plus the fraction x plus 1 over 2 x plus the fraction 5 over x ?

OpenStudy (johnkim2330):

qill give medal @gypsy1274

OpenStudy (johnkim2330):

plz help any1!

OpenStudy (shiraz14):

3x/x = 3 -> 4 + [(x+1)/2x] + (5/x) = (8x + x + 1 + 10)/2x = (9x + 11)/2x = 4.5 + (11/2x)

OpenStudy (johnkim2330):

wait @shiraz14 the choices are 2(x2 + 1) 2x + 1 2x(x + 1) 2(x + 1)

OpenStudy (shiraz14):

The least common denominator should be 2x

OpenStudy (johnkim2330):

\[ \frac{ 3x }{ x+1 } + \frac{ x+1 }{ 2x } + \frac{ 5 }{ x }\]

OpenStudy (shiraz14):

Or wait ... I think I must have misread your question - did you mean 3x/(x+1)? In that case, the least common denominator is 2x(x+1).

OpenStudy (johnkim2330):

oh ok thx @shiraz14

OpenStudy (shiraz14):

You're welcome ... a medal would be much appreciated though ... ;)